OpenStudy (anonymous):

a motor boat took 5h to travel a distance of 60km up a river, against the current. The return trip took 3h. Find the average speed of the boat in still water and the speed of the current

OpenStudy (kropot72):

Let a = average speed of boat in still water Let c = speed of current Upstream speed is given by: \[a-c=\frac{60}{5}\ ................(1)\] Downstream speed is given by: \[a+c=\frac{60}{3}\ ............(2)\] When equations (1) and (2) are added c is eliminated and you can solve to find a the average speed of the boat in still water. Then you can substitute the value of a into equation (1) or equation (2) to find c the speed of the current.